OpenStudy (anonymous):

Use the work-energy theorem to solve. Neglect air resistance in all cases. (a) A branch falls from the top of a 76.0-m-tall redwood tree, starting from rest. How fast is it moving when it reaches the ground? (b) A volcano ejects a boulder directly upward 530 m into the air. How fast was the boulder moving just as it left the volcano? (c) A skier moving at 4.50 m/s encounters a long, rough horizontal patch of snow having coefficient of kinetic friction 0.220 with her skis. How far does she travel on this patch

OpenStudy (shamim):

Work energy theorem is W=k2-k1

OpenStudy (shamim):

W=mgh=m*9.8*76

OpenStudy (shamim):

Kinetic energy at 76 m height is k1=0

OpenStudy (shamim):

Kinetic energy at just above the ground is K2=(1/2)*mv^2

OpenStudy (shamim):

So work energy theorem becomes m*9.8*76=(1/2)*m*v^2 v=?

OpenStudy (shamim):

I wanna help u when u will b online. b,c question

OpenStudy (anonymous):

hello sorry I went to sleep

OpenStudy (anonymous):

How do i solve that if i dont know what m is?

OpenStudy (shamim):

U hv to cancel m frm both side

OpenStudy (shamim):

Feel free to ask for more clarification

OpenStudy (shamim):

After cancellation m frm both side we can write 9.8*76=(1/2)*v^2

OpenStudy (shamim):

v=?